A cement block accidentally falls from rest from the ledge of a 64.7-m-high building. When the block is 18.2 m above the ground, a man, 2.00 m tall, looks up and notices that the block is directly above him. How much time, at most, does the man have to get out of the way?
Answer #1

V^2 = u^2 + 2as

V^2 = 0 + 2 x 9.8 x (64.7 - 18.2)

V = 30.19 m/s

S = ut + 0.5 at^2

-16.2 = - 30.19t - 0.5 x 9.8 x t^2

4.9t^2 + 30.19t - 16.2 = 0

Time, t = 0.50 sec

Comment in case any doubt please rate my answer . .